# Disk Wind Feedback from High-mass Protostars. V. Application of Multi-Modal Machine Learning to Characterize Outflow Properties

### v1

@@ III.2 Interpretable Analysis of Physical Feature Contributions in the Multi-Modal Network
Figure 8: Grad-CAM++ visualizations for outflow mass prediction using the ViT_L_16 model. The background shows the corresponding test outflow input, image (left) and spectrum (right). The overlaid heatmaps indicate regions that contribute most strongly to the inferred quantities.
@@ III.2.1 Smooth Grad-CAM++ Visualization
The resulting heatmaps highlight physically meaningful structures, including central outflow regions, cavity walls, and terminal lobes—that drive the model’s inference. Figure 8 displays representative Grad-CAM++ visualizations for outflow mass prediction using the ViT_L_16 model. A comprehensive comparison, including uncertainty maps and corresponding results for ResNet50, is available in Figure 17 of Appendix B.
We find that both architectures primarily attend to the central outflow and cavity regions when predicting mass, consistent with the expectation that mass is encoded in the overall spatial extent and intensity distribution of the flow. However, distinct attention strategies emerge for position angle prediction (specifically the cosine term), as detailed in Appendix B (Figure 18). While the ResNet50 model focuses sharply on the edges and distal regions of the outflow cavities, ViT_L_16 assigns importance to a broader spatial context, including the void regions surrounding the outflow. This behavior is physically intuitive, as the outflow axis is defined as much by the geometry of the surrounding low-emission regions as by the emitting material itself.
@@ III.2.2 Integrated Gradients with Mean Baseline
Figure 9 displays representative Integrated Gradients maps for mass prediction using the ViT_L_16 model. For a complete comparison, including ResNet50 results, position angle predictions, and associated uncertainty maps, refer to Figures 19 and 20 in Appendix B. The analysis confirms that for both architectures, the dominant contributions stem from central outflow regions and cavity edges. While consistent with the Grad-CAM++ results, IG offers clearer modality separation, revealing that attributions associated with spectral input are systematically weaker than those derived from the images. This contrast is most pronounced in position angle predictions, where the spectral contribution is minimal, and nearly negligible for ViT_L_16. This behavior aligns with physical expectations, confirming that the model correctly isolates spatial morphology as the primary driver for inferring orientation on the plane of the sky.
@@ III.2.3 Occlusion Sensitivity Analysis
Figure 10 presents representative occlusion sensitivity results for mass prediction using the ViT_L_16 model. For a comprehensive comparison, including results for position angle and uncertainty estimates across both ResNet50 and ViT_L_16 architectures, refer to Figures 21 and 22 in Appendix B. The analysis reveals that both architectures exhibit strong sensitivity to the masking of central outflow regions, cavity edges, and terminal lobes, confirming the causal importance of these structures for accurate inference. In contrast, occluding spectral segments produces comparatively minor changes in the model output, providing further independent support for the conclusions drawn from the Integrated Gradients analysis.
@@ IV.2 Comparison with Previous Estimates
The primary source of discrepancy is that our training simulations assume a fixed initial core mass of 60 $M_{\odot}$. As a result, the predicted “protostellar mass” from our models should not be interpreted as an absolute physical mass for real systems, but rather as a proxy for evolutionary stage within the restricted simulation parameter space. Under this interpretation, the ML-inferred masses for G35.20 and G339.88 are broadly consistent with the evolutionary stages inferred from SED modeling. In contrast, G45.47 appears more evolved than suggested by SED fitting when assessed by the stellar-to–initial-core mass ratio, but less evolved in terms of absolute stellar mass, because its initial core mass inferred from SED modeling (228-444 $M_{\mathord{\odot}}$; R. Fedriani et al. 2023) is substantially larger than the fixed initial core mass adopted in our simulations.
@@ V Conclusions
    Across all four backbone architectures evaluated (ResNet50, ResNet152, ViT_B_16, and ViT_L_16), the framework accurately recovers protostellar mass, inclination angle, and position angle. When trained on the full parameter space, all models exhibit low systematic bias and well-constrained uncertainties.
@@
Overall, this work demonstrates that multi-modal learning provides a physically interpretable and scalable approach for linking observed outflow signatures to underlying protostellar properties. While the current model is bounded by the physical diversity of the training simulations, the framework is modular and can be extended to broader parameter spaces, additional molecular tracers, and more heterogeneous environments. These developments will facilitate the automated, physically informed analysis of large-scale ALMA surveys and future high-resolution facilities.
@@
This paper is dedicated to Jan E. Staff (29th May 1977 - 16th May 2023). D.X. acknowledges support from the Virginia Initiative on Cosmic Origins (VICO) and the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number 568580]. D.X. also acknowledges support from the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship Program, a program of Schmidt Sciences. I.A.S. acknowledges support from a Chalmers Astrophysics and Space Sciences Summer (CASSUM) research fellowship. J.C.T. acknowledges support from NSF grant AST-2206450 and ERC Advanced grant 788829 (MSTAR). J.C.T also acknowledges funding from the Virginia Institute for Theoretical Astrophysics (VITA), supported by the College and Graduate School of Arts and Sciences at the University of Virginia.
Appendix A Comparison of Performance for the Four Models
@@
Appendix B Gallery of Model Interpretability Visualizations
@@
In this section, we present a comprehensive gallery of interpretability visualizations comparing the decision-making processes of the ResNet50 and ViT_L_16 architectures. We display Smooth Grad-CAM++ attribution maps for predicted outflow mass and the cosine component of the position angle, along with their respective uncertainty estimates, in Figures 17 and 18. Complementing these, we provide pixel-level attribution via Integrated Gradients for the same parameters in Figures 19 and 20. Finally, the results of the Occlusion Sensitivity Analysis are presented in Figures 21 and 22, verifying the causal importance of the identified spatial features.

### v2

@@ III.2 Interpretable Analysis of Physical Feature Contributions in the Multi-Modal Network
Figure 8: Grad-CAM++ visualizations for outflow mass prediction using the ViT_L_16 model. The background shows the corresponding test outflow input, image (left) and spectrum (right). The overlaid heatmaps indicate regions that contribute most strongly to the inferred quantities. In the left panel, the spatial scale is indicated by a white scale bar representing 0.05 pc. In the right panel, the x-axis displays the line-of-sight velocity (in km s-1), and the y-axis represents the normalized intensity (scaled from 0 to 1).
@@ III.2.1 Smooth Grad-CAM++ Visualization
The resulting heatmaps highlight physically meaningful structures that drive the model’s inference. These include the central outflow regions (the dense, high-velocity gas immediately surrounding the protostellar launch site), the cavity walls (the swept-up material delineating the boundary between the outflowing gas and the ambient envelope), and the terminal lobes (the extended, shocked structures at the leading edge where the outflow impacts the surrounding medium). These synthetic features correspond directly to structures commonly seen in real protostellar outflows. For instance, well-defined cavity walls and terminal bow shocks are prominently observed in high-resolution studies of low-mass systems such as HH 212 (C.-F. Lee et al., 2021; J. A. López-Vázquez et al., 2024) and L1527 (N. Sakai et al., 2017), as well as in high-mass counterparts like Orion-KL Source I (e.g., M. Wright et al., 2023) and the ALMA targets analyzed in Section IV. Figure 8 displays representative Grad-CAM++ visualizations for protostellar mass prediction using the ViT_L_16 model. A comprehensive comparison, including uncertainty maps and corresponding results for ResNet50, is available in Figure 20 of Appendix C.
We find that both architectures primarily attend to the central outflow and cavity regions when predicting mass, consistent with the expectation that mass is encoded in the overall spatial extent and intensity distribution of the flow. However, distinct attention strategies emerge for position angle prediction (specifically the cosine term), as detailed in Appendix C (Figure 21). While the ResNet50 model focuses sharply on the edges and distal regions of the outflow cavities, ViT_L_16 assigns importance to a broader spatial context, including the void regions surrounding the outflow. This behavior is physically intuitive, as the outflow axis is defined as much by the geometry of the surrounding low-emission regions as by the emitting material itself.
@@ III.2.2 Integrated Gradients with Mean Baseline
Figure 9 displays representative Integrated Gradients maps for mass prediction using the ViT_L_16 model. For a complete comparison, including ResNet50 results, position angle predictions, and associated uncertainty maps, refer to Figures 22 and 23 in Appendix C. The analysis confirms that for both architectures, the dominant contributions stem from central outflow regions and cavity edges. While consistent with the Grad-CAM++ results, IG offers clearer modality separation, revealing that attributions associated with spectral input are systematically weaker than those derived from the images. This contrast is most pronounced in position angle predictions, where the spectral contribution is minimal, and nearly negligible for ViT_L_16. This behavior aligns with physical expectations, confirming that the model correctly isolates spatial morphology as the primary driver for inferring orientation on the plane of the sky.
@@ III.2.3 Occlusion Sensitivity Analysis
Figure 10 presents representative occlusion sensitivity results for mass prediction using the ViT_L_16 model. For a comprehensive comparison, including results for position angle and uncertainty estimates across both ResNet50 and ViT_L_16 architectures, refer to Figures 24 and 25 in Appendix C. The analysis reveals that both architectures exhibit strong sensitivity to the masking of central outflow regions, cavity edges, and terminal lobes, confirming the causal importance of these structures for accurate inference. In contrast, occluding spectral segments produces comparatively minor changes in the model output, providing further independent support for the conclusions drawn from the Integrated Gradients analysis.
@@ IV Application to Real ALMA-Observed Protostellar Outflows and Uncertainty Assessment
Figure 12: Side-by-side comparison of real ALMA 12CO (2-1) observations (top row) and synthetic counterparts (bottom row) retrieved from the simulation grid using the model’s predicted mass, inclination, and position angle. While stochastic turbulent features prevent a pixel-perfect match, the global morphology and cavity geometry show strong qualitative agreement.
@@ IV.2 Comparison with Previous Estimates
The primary source of discrepancy is that our training simulations assume a fixed initial core mass of 60 $M_{\odot}$. As a result, the predicted “protostellar mass” ($m_{*}$) from our models should not be interpreted as an absolute physical mass for real systems, but rather as a proxy for the evolutionary fraction of the core ($m_{*}/M_{\rm core}$) within the restricted simulation parameter space. Our interpretability analysis supports this: the feature importance maps indicate that the network relies heavily on the spatial morphology of the outflow (e.g., the widening of the cavity walls) to infer the mass. However, we note that the outflow opening angle is not strictly a dimensionless geometric function of $m_{*}/M_{\rm core}$. Cavity widening is fundamentally a momentum-driven process that depends strongly on the wind mass-loss rate and, by extension, the absolute mass accretion rate (Y. Zhang et al., 2014). Consequently, the model effectively learns the specific evolutionary sequence—and the associated momentum-injection history—of a massive 60 $M_{\odot}$ core . This single evolutionary track introduces an inherent degeneracy between the protostellar mass and the mass accretion rate, meaning the inference may be biased when applied to protostars embedded in significantly lower- or higher-mass envelopes with different accretion histories. Under this interpretation, the ML-inferred masses for G35.20 and G339.88 are broadly consistent with the evolutionary stages inferred from SED modeling. In contrast, G45.47 appears more evolved than suggested by SED fitting when assessed by the stellar-to-initial-core mass ratio, but less evolved in terms of absolute stellar mass, because its initial core mass inferred from SED modeling (228-444 $M_{\mathord{\odot}}$; R. Fedriani et al. 2023) is substantially larger than the fixed initial core mass adopted in our simulations.
@@
To provide a direct visual validation of these results, we present Figure 12, which provides a side-by-side comparison between the real ALMA 12CO (2–1) observations of our three targets and synthetic counterparts retrieved from the simulation grid using our predicted physical parameters. While some real observed outflows, such as G35.20, may only display a single prominent lobe due to environmental complexities or source multiplicity, the current machine learning model successfully identifies the underlying physical structure by finding the closest global match for mass and inclination. The qualitative agreement in lobe shape and collimation demonstrates that the model effectively maps observed morphologies back to the underlying physical properties of the disk-wind system.
This qualitative approach to validation is necessitated by the nature of the 3D MHD simulations, in which the envelopes and outflows undergo stochastic structural evolution. Because these simulations represent specific physical realizations of a turbulent core rather than deterministic, pixel-level templates, the detailed structure of features such as cavity walls or local intensity fluctuations is not expected to match real-world sources exactly. Instead, as demonstrated in our interpretability analysis, the multi-modal framework drives its inference from robust, global morphological markers—such as cavity geometry and lobe extent—that are resolved in the integrated intensity maps. The visual consistency in these large-scale features confirms that the model is capturing the fundamental physical drivers of the outflow rather than overfitting to stochastic substructures.
@@ IV.3 Position Angle as a Robust Validation Metric
Beyond the physical parameters of the training set, a primary consideration for broader application is the potential domain mismatch when transitioning between observational facilities or molecular tracers. Our model is currently optimized for the specific spatial frequency and noise characteristics of ALMA interferometric data, which differ significantly from the continuous spatial sensitivity of single-dish telescopes or unfiltered simulation outputs. Furthermore, the framework is strictly calibrated for the 12CO (2-1) transition based on the specific excitation and abundance profiles of our MHD training simulations. Tracers such as SiO, SO2, or HCO+ involve distinct chemical and excitation physics, including shock-induced enhancement, selective depletion, and significantly higher critical densities, that are not captured in the current model weights. Because these species often trace higher-density gas or specific shock conditions, the mapping between their observed intensity and global parameters like mass or inclination would differ from that of the more ubiquitous CO emission. Consequently, direct application to different instruments or tracers could result in performance degradation due to these differing physical and instrumental responses. To ensure high fidelity across diverse setups, a “fine-tuning” or transfer learning step would be required. This involves re-calibrating the pre-trained model using a specialized synthetic dataset processed through the target instrument’s specific response function or the appropriate radiative transfer physics of a new molecular tracer, allowing the network to adapt to new conditions while preserving the fundamental physical correlations learned from the underlying simulations.
@@ V Conclusions
    Across all four backbone architectures evaluated (ResNet50, ResNet152, ViT_B_16, and ViT_L_16), the framework accurately recovers protostellar properties; specifically, when trained on the full parameter space, all models exhibit low systematic bias, whereas tests on restricted datasets reveal that Transformer-based architectures are more robust to morphological degeneracies than ResNet models.
@@
Overall, this work demonstrates that multi-modal learning provides a physically interpretable and scalable approach for linking observed outflow signatures to underlying protostellar properties. However, it is vital to explicitly acknowledge the physical limitations inherent to the current training dataset. Most notably, because the underlying MHD simulations utilize a fixed initial core mass of $60~M_{\odot}$, the model is trained on a single evolutionary track. Consequently, the predicted protostellar mass ($m_{*}$) is degenerate with the mass accretion rate and primarily acts as a proxy for the evolutionary stage ($m_{*}/M_{\rm core}$) of a massive core, rather than representing a universal absolute mass. Furthermore, the current framework is trained exclusively on synthetic 12CO (2-1) emission. While bounded by these physical and observational constraints, the framework is modular. Future expansions of the simulation grid to include varied initial core masses and envelope densities, combined with targeted transfer learning for additional molecular tracers, will break these physical degeneracies. Ultimately, these developments will facilitate the automated, physically informed analysis of large-scale ALMA surveys and future high-resolution facilities.
@@
This paper is dedicated to Jan E. Staff (29th May 1977 - 16th May 2023). We thank the anonymous referee for their careful reading of the manuscript and for their constructive and insightful comments. Their suggestions have significantly improved the clarity and robustness of this work. D.X. acknowledges support from the Virginia Initiative on Cosmic Origins (VICO) and the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number 568580]. D.X. also acknowledges support from the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship Program, a program of Schmidt Sciences. I.A.S. acknowledges support from a Chalmers Astrophysics and Space Sciences Summer (CASSUM) research fellowship. J.C.T. acknowledges support from NSF grant AST-2206450 and ERC Advanced grant 788829 (MSTAR). J.C.T also acknowledges funding from the Virginia Institute for Theoretical Astrophysics (VITA), supported by the College and Graduate School of Arts and Sciences at the University of Virginia. Computations were performed on the Trillium supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto
Appendix A Training Dynamics and Learning Curves
To provide transparency regarding the model training process and to demonstrate the convergence stability of our multi-modal framework, we present the learning curves for both the ViT_L_16 and ResNet50 architectures. Figures 14 and 15 display the evolution of the Gaussian Negative Log-Likelihood (NLL) for both the training and validation sets, alongside the validation Mean Squared Error (MSE) plotted on a logarithmic scale.
As detailed in Section II.3, our training procedure utilizes a progressive fine-tuning strategy combined with heavy, on-the-fly data augmentation. During the initial training phases, the models rapidly optimize the NLL loss on the training data. Because the NLL loss function evaluates both the predicted mean and the predicted log-variance ($s=\ln(\sigma^{2})$), the network can initially become overly confident, predicting artificially narrow variances for the training set. When applied to the unseen validation set, this overconfidence temporarily drives up the validation NLL, despite the underlying validation MSE remaining relatively stable.
However, at the scheduled learning rate decay and layer unfreezing milestones, the network successfully recalibrates its uncertainty estimations. This recalibration is visually evident in the learning curves as sudden, sharp drops in the validation NLL, accompanied by corresponding decreases in the validation MSE. These drops indicate that the models have escaped local minima and successfully generalized to the validation data.
Our early-stopping protocol is designed to capture the model weights exactly at these global minima, halting the training process before long-term overfitting can degrade performance. For the fiducial ViT_L_16 framework (Figure 14), the lowest validation NLL was achieved at Epoch 392, yielding a validation NLL of 4.93-4.93, an associated training NLL of 4.88-4.88, and a validation MSE of $7\times 10^{-6}$. Similarly, the ResNet50 framework (Figure 15) reached its optimal performance at Epoch 4273, with a validation NLL of 4.61-4.61, a training NLL of 4.25-4.25, and a validation MSE of $9\times 10^{-6}$. The tight alignment between the training and validation NLL at these specific stopping epochs confirms that the networks are well-calibrated and have learned robust physical representations rather than memorizing the discrete simulation grid.
Figure 14: Learning curve for the ViT_L_16 multi-modal model. The plot shows the training NLL (blue solid line), validation NLL (red solid line), and validation MSE (green dashed line, logarithmic scale on the right axis) as a function of training epoch. The sharp drop near epoch 350 corresponds to a scheduled learning rate decay and layer unfreezing, leading to the optimal model weights saved at Epoch 392.
Figure 15: Learning curve for the ResNet50 multi-modal model. The plot shows the training NLL (blue solid line), validation NLL (red solid line), and validation MSE (green dashed line, logarithmic scale on the right axis). The model successfully recalibrates after progressive fine-tuning milestones, achieving its lowest validation loss at Epoch 4273.
Appendix B Comparison of Performance for the Four Models
@@
Appendix C Gallery of Model Interpretability Visualizations
@@
In this section, we present a comprehensive gallery of interpretability visualizations comparing the decision-making processes of the ResNet50 and ViT_L_16 architectures. We display Smooth Grad-CAM++ attribution maps for predicted outflow mass and the cosine component of the position angle, along with their respective uncertainty estimates, in Figures 20 and 21. Complementing these, we provide pixel-level attribution via Integrated Gradients for the same parameters in Figures 22 and 23. Finally, the results of the Occlusion Sensitivity Analysis are presented in Figures 24 and 25, verifying the causal importance of the identified spatial features.